OpenStudy (mathhelp346):

The coin box of a vending machine contains 6 times as many quarters as dimes. If the total amount of money in quarters and dimes is $28.80, how many quarters and how many dimes are in the box? Use the BOX METHOD please!

OpenStudy (anonymous):

Let no of dimes =10x Then no of qtrs=60x value of dimes=10x/10=1x value of qtrs=60x/4=15x total=15x+1x=16x=28.8 $ 16x=28.8 x=28.8/16=7.2/4=1.8

OpenStudy (anonymous):

no of dimes=1.8x10=18 vo of qtrs=18x6=108 verification: value of 108 qtrs=108/4=27 $ value of 18 dimes=1.8 $ total=27+1.8=28.8 $

OpenStudy (mathhelp346):

why is no of quarters 60x

OpenStudy (anonymous):

because a vending machine contains 6 times

OpenStudy (anonymous):

10 * 6 =60

OpenStudy (mathhelp346):

so x is quarters right

OpenStudy (mathhelp346):

but wouldnt it be 6x?

OpenStudy (mathhelp346):

so in the box method would the 2 equations be 25q+10d=2880 60q+10d=x

OpenStudy (anonymous):

Dude whatever steps I showed you above just follow it or ask someone else because I explained what I had to.